OpenStudy (anonymous):

Solve using the square root method. Root the solution to the nearest tenth. x^2 - 20= 0

OpenStudy (anonymous):

\[x^2 - 20 = 0\] \[(x + \sqrt{20})(x - \sqrt{20}) = 0\] \[x = \sqrt{20} \quad Or \quad x = -\sqrt{20}\]

OpenStudy (anonymous):

Find the square root of 20 and you will get your answer..

OpenStudy (anonymous):

thanks :)